A uniform piece of wire has a diameter of 0.30 mm and length 2.25 m has a resistance of 28 Ω. Find the resistivity of the wire. 


1
Expert's answer
2022-02-25T11:11:57-0500

The resistivity of a conductor is given by

"R=\rho \dfrac{L}{A}"

Where 

"A" is the Area

"\rho" is the Resistivity

"L" the length



the radius is 

"r=\dfrac{d}{2} \\
r=\dfrac{0.30mm}{2}\\
r=0.15mm"

Converting from millimeters to meters

"r=0.15mm\times\dfrac{1m}{1000mm}=0.00015m"


The wire area is

"A=\pi r^{2}"

Evaluatin numerically

"A=\pi (0.00015m)^{2}"

"A=7.07\times10^{-8}m^{2}"

Getting the expression for the resistivity

"R=\rho \dfrac{L}{A}"

"\dfrac{A}{L}R=\rho"


"\rho=\dfrac{A}{L}R"

Evaluating numerically

"\rho=\dfrac{7.07\times10^{-8}m^{2}}{2.25m}\times 28 \Omega"



The resistivity is

"\rho=8.80\times10^{-7}\Omega m"